OpenStudy (anonymous):

lim x→infinity x^5/x

OpenStudy (zzr0ck3r):

= lim x->infinity x^4

OpenStudy (anonymous):

but x is infinity...

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}x^4= \infty\]

Parth (parthkohli):

\[\rm \lim_{x \to \infty} {x^5 \over x} = \lim_{x \to \infty} x^4 = \infty\]

OpenStudy (anonymous):

nope, is no infinity . i tried but i cant get the right answer....